Find the regression​ equation, letting the first variable be the predictor​ (x) variable. Using the listed​lemon/crash data, where lemon imports are in metric tons and the fatality rates are per​ 100,000 people, find the best predicted crash fatality rate for a year in which there are 425 metric tons of lemon imports. Is the prediction​ worthwhile?

Lemon Imports

228

263

350

482

525

Crash Fatality Rate

15.9

15.7

15.5

15.3

14.8

A. Find the equation of the regression line.

^y=_+(_)x

B. The best predicted crash fatality rate for a year in which there are 425 metric tones of lemon imports is ___ fatalities per 100,000 population.

C. Is the prediction worthwhile?

-a. Since the sample size is​ small, the prediction is not appropriate.

-b. Since common sense suggests there should not be much of a relationship between the two​variables, the prediction does not make much sense.

-c. Since all of the requirements for finding the equation of the regression line are​ met, the prediction is worthwhile.
Answer #1

(y-y) 0.2116 0.0676 0.0036 0.0196 0.4096 0.7120 (x-x)(y-y 65.1360 27.7160 1.1760 -15.7360 99.4560 209.2200 S.No (AT) X-X 263 350 482 525 1848 369.6000 15.9 15.7 15.5 15.3 14.8 77.2 15.4400 20050.5600 11363.5600 384.1600 12633.7600 24149.1600 68581.2000 2 4 Total Mean SSY sample size here 15.4400 369.6000 from above 68581.2000 209.2200 slope- -0.0031 intercept- 16.5675 hence line equation: 16.5675 0.0031 x

a)

frm above line equation: Yhat=16.568+(-0.003)*x

b)

predicted crash fatality =16.568+(-0.003)*425=15.29 ( please try 15.27 if this comes wrong due to rounding error)

c)

b. Since common sense suggests there should not be much of a relationship between the two​variables, the prediction does not make much sense.